Variable Intercept A study of 35 secretaries yearly salaries (in thousands of dollars) was done. The researchers want to preIdentify the Test Statistic and P-Value
Answer #1

c) H0 : Typing speed contributed nothing useful after allowing for the other predictors in the model, B4 = 0

HA : Typing speed makes a useful contribution to the model, 07 g

The value of test statistic t = 0.026

P-value

= P(n-6 > tobs)

= P(|t29 > 0.026)

= 2* P(t29 > 0.026)

= 2* [1 - P(t29 < 0.026)

=2+1-0.5102823)

=2+0.4897177

= 0.9794354

<0.9794